I have to find the distance between (-9,4)and(0,-8) can i solve it this way?
4+-8=4
-9+0=-9

Answer by praseenakos@yahoo.com(507)

QUESTION:

find the distance between (-9,4)and(0,-8)


ANSWER:


Here you have to use distance formula to find the distance between two points.
It is given that,

Distance = sqrt of {(difference of x coordinates)^2 + (difference of y coordinates)}


(Difference of x coordinates)^2 = (-9 - 0)^2 = (-9)^2 = 81

(Difference of y coordinates )^2 = ( 4 - -8)^2 = (4 + 8 ) ^2 = 12^2 = 144

So Distance = sqrt of (81 + 144)

==> = sqrt of (125)



==> = 15


So distance between the given points is 15.



Hope you find the explanation useful.



Regards.


Praseena.